# Unsupervised Cross-Domain 3D Human Pose Estimation via Pseudo-Label-Guided Global Transforms

### v1

@@ III-A Problem Formulation
III-B Overview of the Proposed Method
Our proposed framework aims to synthesize new 2D-3D pose pairs from the labeled source domain and the unlabeled target domain, and then use them to train a pose estimator to predict the 3D human pose in the target domain. Among the various factors contributing to the domain gap, camera viewpoints and positions play a significant role by affecting the global positions of the poses. Thus, the key challenge to be met is how to make the augmented data adequately match the distribution of the target dataset. To achieve this, we explore the latent knowledge of the target dataset as follows (also illustrated in Fig. 2). A Pseudo-Label Generation module $M_{p}$ for the target dataset is proposed to generate pseudo-3D poses. For the source dataset, we employ an augmentation module, Pose Augmentor $M_{a}$, that creates more diverse samples to better account for motions and body sizes. The 2D and 3D poses from both domains are then spatially aligned in a Global Transformation module $M_{t}$ using a human-centric coordinate system as a bridge. These aligned 2D-3D pairs are then fed into a Pose Discriminator $M_{d}$ to distinguish reliable poses to be used to fine-tune the Pose Estimator $M_{e}$. Finally, the Pose Estimator $M_{e}$ is applied to the target dataset to generate updated 3D pseudo-labels. Since the Global Transformation module is guided by pseudo-labels, updated pseudo-labels enhance alignment between synthesized and target data. Therefore, it may be possible to boost the pose estimation results by iterating the above process. We explore this in our ablations.
III-C Pseudo-Label Generation
For the target dataset, we generate 3D pseudo-labels by inferring absolute 3D coordinates of human skeleton joints in the camera coordinate system from the available 2D labels. These 3D pseudo-labels can be obtained using a pre-trained pose estimation method given 2D inputs.
Given a 2D pose $\bm{P}_{2D}^{tgt}=[\bm{x}^{tgt},\bm{y}^{tgt}]^{T}$ from the target dataset, we obtain the root-relative 3D coordinates $\hat{\bm{P}}_{3D}^{tgt}=[\hat{\bm{X}}^{tgt},\hat{\bm{Y}}^{tgt},\hat{\bm{Z}}^{tgt}]^{T}$ using VideoPose3D [23], pretrained on the source dataset [^1] We tried other methods, e.g. MixSTE [24], obtaining very similar results. The predicted 3D coordinates of the joints $(\hat{X_{j}}^{tgt},\hat{Y_{j}}^{tgt},\hat{Z_{j}}^{tgt}),j=\left\{0,...,J-1\right\}$ are root-relative 3D coordinates with the root $(\hat{X_{0}}^{tgt},\hat{Y_{0}}^{tgt},\hat{Z_{0}}^{tgt})$ set to $(0,0,0)$. To obtain the absolute coordinates of the root, we apply a two-stage optimization process. In the first stage, given the 2D pose and camera intrinsic parameters $\bm{K}$, the 3D pose can be inferred by assuming an initial depth value.
The corresponding bone lengths are then derived from the estimated 3D pose. To ensure anatomical consistency, we minimize the discrepancy between the estimated bone lengths and their averaged values using least squares, and thus, we obtain the 3D absolute coordinates of all joints $\tilde{\bm{P}}_{j}=({\tilde{X}_{j}}^{tgt},{\tilde{Y}_{j}}^{tgt},{\tilde{Z}_{j}}^{tgt}),j=\left\{0,...,J-1\right\}$. In the second stage, we minimize the error between reprojected 2D pose from $\hat{\bm{P}}_{3D}^{tgt}+\bm{P}_{r}$ and real 2D pose $\bm{P}_{2D}^{tgt}$, where $\bm{P}_{r}$ is the absolute coordinates of the root joint and its initial value is set to root joint $\tilde{\bm{P}}_{0}$. Finally, we arrive at the estimated absolute 3D coordinates of all joints $\tilde{\bm{P}}_{3D}^{tgt}$ as the 3D pseudo-labels of the target dataset.
III-D Pose Augmentor
The poses in the source dataset are augmented using the bone augmentation process from AdaptPose [18]. This allows our Pose Augmentor $M_{a}$ to modify bone angles across frames to introduce greater diversity in motion representation and ensure the augmented poses capture a wider range of actions. Additionally, bone lengths are adjusted to accommodate variations in human body proportions, enhancing the model’s adaptability to different body sizes.
Given a sequence of 3D poses $\left\{\bm{P}_{3D}^{src}\right\}_{n=0}^{N-1}$ from the source dataset, our augmentor converts them into their corresponding bone vectors $\left\{\bm{B}^{src}\right\}_{n=0}^{N-1}\in R^{N\times(J-1)\times 3}$. Each original bone vector $b^{l}_{n}\in R^{1\times 3}$ (i.e., the $l$-th bone vector from $\bm{B}^{src}$ in $n$-th frame) is rotated about an axis $\bm{w}\in R^{(J-1)\times 3}$ by a certain rotation angle $\theta\in R^{(J-1)\times 3}$ proportional to the number of frames. This guarantees that the bone changes are temporally plausible and consistent. Additionally, to adjust the bone length, the pose augmentor scales it using a ratio parameter $\lambda\in R^{(J-1)\times 3}$, which systematically modifies the bone length while preserving the structural coherence of the motion.
To predict the parameters ($\bm{w},\theta,\lambda$), a Gaussian-noise vector is concatenated with the input 3D pose of the middle frame to introduce randomness and diversity for feature extraction. Subsequently, four fully connected layers followed by batch normalization and a leaky ReLU are used to extract features and learn these parameters [18].
@@ III-E Global Transformation
The Global Transformation module $M_{t}$ is designed to eliminate differences in camera viewpoints and positions between the source and target domains. To achieve the optimal alignment of 3D coordinates from both domains, we propose a human-centric coordinate system as a bridge. As shown in Fig. 3, the origin of this coordinate system $H$ is centered at the middle of the left and right hip joints. The x-axis is defined as the unit vector pointing from the origin to the right hip, while the x-y plane is defined by three joints: the left hip, right hip, and thorax joint. The z-axis is determined by the unit vector perpendicular to x-y plane and y-axis is the cross-product of x and z axes.
Given an augmented 3D pose $\bm{P}_{3D}^{\mathcal{A}}$ from the source dataset under the source camera coordinate system $C_{s}$, we transform it using rotation and translation matrices ($\bm{R}_{1},\bm{T}_{1}$) from $C_{s}$ to $H$, where $H$ is our human coordinate system (see Fig. 3). Similarly, given a 3D pseudo-pose $\tilde{\bm{P}}_{3D}^{tgt}$ sampled from the target dataset (that were generated earlier in Section III-C) under the target camera coordinate system $C_{t}$, we compute the corresponding rotation and translation matrices ($\bm{R}_{2},\bm{T}_{2}$) from $C_{t}$ to $H$. The transformation from any camera coordinate system to the human-centric coordinate system $H$ is outlined in Algorithm 1 (See Appendix).
Using the human coordinate system $H$ as a bridge, the discrepancy between the source and target domains in terms of positions can be determined from variations of the subject’s position and orientation relative to the camera. Thus, to eliminate this discrepancy, we first obtain
$\bm{P}_{H}=\bm{R}_{1}\cdot(\bm{P}_{3D}^{\mathcal{A}}+\bm{T}_{1}),$ (3)
where $\bm{P}_{H}$ denotes the 3D human pose in the human centered coordinate system. Then, we convert $\bm{P}_{H}$ into the target camera coordinate system $C_{t}$ to enable the synthesis of new data from $\bm{P}_{3D}^{\mathcal{A}}$, such that
$\bm{P}_{3D}^{{}^{\prime}}=\bm{R}_{2}^{-1}\cdot\bm{P}_{H}-\bm{T}_{2},$ (4)
@@
where ${\bm{P}}_{3D}^{{}^{\prime}}=[\bm{X}^{{}^{\prime}},\bm{Y}^{{}^{\prime}},\bm{Z}^{{}^{\prime}}]^{T}\in R^{3\times J}$, ensuring alignment with the target domain. Finally, we can project ${\bm{P}}_{3D}^{{}^{\prime}}$ into 2D pose ${\bm{P}}_{2D}^{{}^{\prime}}=[\bm{u},\bm{v}]^{T}\in R^{2\times J}$ where $(\bm{u},\bm{v})$ denote the 2D image coordinates, obtained as
$\begin{array}[]{l}u_{j}=f_{x}\cdot\frac{X_{j}^{{}^{\prime}}}{Z_{j}^{{}^{\prime}}}+c_{x},\\[10.0pt] v_{j}=f_{y}\cdot\frac{Y_{j}^{{}^{\prime}}}{Z_{j}^{{}^{\prime}}}+c_{y},~{}j=\left\{0,...,J-1\right\}\end{array}$ (5)
@@
where $(f_{x},f_{y},c_{x},c_{y})$ are the target dataset camera’s intrinsic parameters. As shown in Fig. 5, after global transformation, 3D poses from the source and target domains are optimally aligned. For 2D poses, their root positions are also aligned and the scales are similar. This alignment process significantly reduces the domain discrepancies caused by differences in camera viewpoints and positions.
@@
Figure 5: Visualization of the global transformation. The target pseudo-3D pose inferred from the target 2D pose guides rotation and translation of the source 3d pose to generate the aligned 3D pose. Given the aligned 3D poses, corresponding 2D poses are obtained through projection using the intrinsic camera parameters of the target dataset so that the 2D root positions are also aligned.
@@ III-F Iterative training architecture
After obtaining the aligned 2D and 3D poses ${\bm{P}}_{2D}^{{}^{\prime}}$ and ${\bm{P}}_{3D}^{{}^{\prime}}$ respectively, we train an adversarial pose augmentation framework to ensure the fidelity of the generated human poses (following previous works [17, 21]). This augmentation framework involves our generator, i.e., the pose augmentor $M_{a}$ and the discriminator $M_{d}$. The discriminator performs two tasks. First, it uses the aligned 2D pose ${\bm{P}}_{2D}^{{}^{\prime}}$ and the target 2D pose ${\bm{P}}_{2D}^{tgt}$ at the middle frame as input, to adjust ${\bm{P}}_{2D}^{{}^{\prime}}$ to closely match the target sample. This is unlike PoseAug [17] which uses the 2D poses of source dataset as input. In the second task, the discriminator uses the middle-frame 3D augmented pose ${\bm{P}}_{3D}^{\mathcal{A}}$ and the source pose ${\bm{P}}_{3D}^{src}$ as input to ensure the fidelity of poses after bone angle and bone length augmentation. Inspired by Kinematic Chain Space (KCS) [49, 50], we use KCS representation of 3D pose as input rather than 3D joints.
In each epoch, the generator $M_{a}$, discriminator $M_{d}$ and the pose estimator $M_{e}$ are trained jointly. Their loss functions are as follows.
@@
Generator loss. The adversarial loss of the generator $M_{a}$ is the feedback from the discriminator:
@@
where $\bm{P}$ and $\tilde{\bm{P}}$ are the ground truth and predicted 3D poses respectively. Before training $M_{e}$, the generator and the discriminator need to be warmed up for $m$ epoch.
After the training process, we obtain a fine-tuned Pose Estimator $M_{e}$ which generates predicted 3D poses $\hat{\bm{P}}_{3D}^{tgt}$ for the target dataset. These poses could be used as input for the Pseudo-Label Generation module $M_{p}$ again, and thus, the training process can be iterative and the number of iterations are studied in the ablation study.
@@ IV-A Datasets and Metrics
Human3.6M (H3.6M) is an extensive dataset for 3D human pose estimation, captured in a controlled indoor environment. The dataset includes over 3.6m frames, covering seven main subjects (S1, S5, S6, S7, S8, S9, and S11) engaged in 15 diverse activities. Following previous works [21, 20], H3.6M has two main settings: (i) for cross-dataset evaluation, we use subjects S1, S5, S6, S7, and S8 as the source domain, with the other dataset as the target domain, and (ii) for cross-scenario evaluation, we use only S1 as the source domain, with S5, S6, S7, and S8 as the target domain. Measures MPJPE and PA-MPJPE are standard metrics employed to evaluate performance across these tasks.
@@ IV-B Implementation Details
Pose Estimator. Following previous works [17, 18, 22], we use single-frame-based VideoPose3D [23] and video-based MixSTE [24] for our Pose Estimator $M_{e}$. The input sequence lengths of MixSTE was set to 27 for 3DHP and 3DPW, and 243 for Human3.6M following the settings in [22]. The pre-trained weights from the source dataset are used as initial weights. We adopt the 16-keypoint human model with 0-th joint (hip center point) as the origin and we use ground truth 2D keypoints as input.
Generators and discriminators. The generator $M_{a}$ has three 3-layer residual MLPs with LeakyReLU to generate augmented bone angles and bone lengths. We set $\alpha=1$ and $\beta=6$ in Eq. 7. The Pose Discriminator $M_{d}$ uses the same architecture for processing both 2D and 3D poses, consisting of five 4-layer residual MLPs with LeakyReLU with each corresponding to a part of KCS representation.
Training. The initial learning rates for the discriminator, generator, and pose estimator are all set to 0.0001. They all use Adam optimizer. The experiments are conducted with a batch size of 1024 for 50 epochs and there is a warm-up phase lasting 2 epochs to train the generator and discriminator.
@@ IV-C Results
In the first, H3.6M is the source and 3DHP iss the target datset (see Table I). For single frame-based VideoPose3D [23] as the Pose Estimator $M_{e}$, our method outperforms PoSynDA[22] by 2.5$\downarrow$ on MPJPE and PoseDA[21] by 2.1$\uparrow$ on AUC. For video-based MixSTE [24] as the Pose Estimator, our method improves on PoSynDA[22] by 4.2$\downarrow$ in MPJPE, 0.9$\uparrow$ in PCK and 5.4$\uparrow$ in AUC. Note, our method using MixSTE [24] surpasses the target-specific, fully supervised MixSTE [24] method itself (from 54.9mm to 54.0mm in MPJPE).
@@
In the second experiment, H3.6M is the source and 3DPW is the target dataset (see Table II). For both single-frame and video-based methods, our approach achieve state-of-the art results. Amongst single-frame methods, ours surpasses PoseDA[21] in MPJPE by 12.0$\downarrow$ and in PA-MPJPE by 10.2$\downarrow$ respectively. For video-based approaches, again in comparison with the state-of-the-art PoSynDA[22], we improve by 8.9$\downarrow$ and 3.5$\downarrow$ in MPJPE and PA-MPJPE, respectively. Our method manages greater strides in improved results on 3DPW than on 3DHP, as it is designed to address the issue of variations in camera position and orientation which are more pertinent to the dynamic camera motion scenes in 3DPW.
@@
$\begin{array}[]{l}\begin{bmatrix}X_{new}\\ Y_{new}\\ Z_{new}\end{bmatrix}=\begin{bmatrix}1&0&0\\ 0&\cos(45^{\circ})&\sin(45^{\circ})\\ 0&\sin(45^{\circ})&\cos(45^{\circ})\end{bmatrix}\cdot\begin{bmatrix}X\\ Y+2\\ Z\end{bmatrix},\\ \begin{bmatrix}u_{new}\\ v_{new}\\ 1\end{bmatrix}=\begin{bmatrix}f_{x}^{p}&0&c_{x}^{p}\\ 0&f_{y}^{p}&c_{y}^{p}\\ 0&0&1\\ \end{bmatrix}\cdot\begin{bmatrix}\frac{X_{new}}{Z_{new}}\\[3.0pt] \frac{Y_{new}}{Z_{new}}\\[3.0pt] 1\end{bmatrix},\end{array}$ (9)
@@
Table III presents comparative results against PoseAug[17] and AdaptPose[18] with H3.6M as source and 3DHP-C as the target dataset. We do not compare with PoseDA [21] as their complete codes are not available, nor with PoSynDA [22] since we obtained excessively incorrect results using their provided code. Our method outperforms by a large margin. This improvement stems from our explicit handling of camera viewpoint discrepancies which impresses the potential effectiveness of the proposed method under challenging viewpoints.
@@
Qualitative evaluation. Fig. 6 (a)-(d) presents example qualitative results on the 3DHP dataset which show that compared to the pre-trained baseline model VideoPose3D[23], our method performs significantly better, especially for joints that are farther away from the root joint. Fig. 6(e) illustrates a failure case of our method where the estimation of joints on the right side of the body is not accurate enough due to extreme self-occlusion.
Figure 6: Example 3D pose estimations on 3DHP dataset. (a)-(d) Our method (blue) performs significantly better than the pre-trained model (green) within smaller error compared with the ground truth (red). (e) In cases of severe self-occlusion, our method fails to obtain sufficiently estimations.

### v2

@@ III-E Global Transformation
Camera viewpoints and positions across datasets and/or domains lead to significant discrepancies in the 3D absolute coordinates of human poses. Given a 3D human pose $P^{A}$ captured from a source domain and another 3D pose $P^{B}$ from a target domain, the objective would be to eliminate the domain discrepancy by transforming $P^{A}$ so that it aligns with $P^{B}$. This problem can be formulated as an optimal rigid alignment task, where the goal is to estimate the rotation matrix $R$ and translation matrix $T$ to minimize $\left\|R\cdot P^{A}+T-P^{B}\right\|$.
A well-established solution to this is the Kabsch algorithm [50, 51], which computes the cross-covariance matrix between the two point sets, followed by Singular Value Decomposition to derive the optimal transformation. However, this process is complex and time-consuming, especially when applied to the training process on large datasets.
Inspired by prior works [52, 53] which introduced a body-centered coordinate system to achieve viewpoint invariance, we propose to use a human-centric coordinate system $H$ as a bridge to align 3D coordinates across domains. A 3D human pose $P$ can be expressed as $P=P_{r}+\Delta P$, where $P_{r}$ denotes the 3D absolute coordinate of the root which usually is the middle hip joint, and $\Delta P$ is the relative coordinates compared to the root. The transformation of $P^{A}$ to align with $P^{B}$ incorporates three steps. The first step is to center $P^{A}$ at $H$ by subtracting the root’s location $P_{r}^{A}$. In the second step, the aim is to find a rotation to minimize the error between joints’ relative coordinates. In the third step, a translation is applied to align the root joint with $P_{r}^{B}$. The theoretical derivations can be found in Appendix A.
@@
As can be seen in Fig. 4, the origin of the human-centric coordinate system is taken as the center of the left and right hip joints (labeled 0-hip). The x-axis is defined as the unit vector pointing from the right hip to the left hip, while the x-y plane is defined by three joints: the right hip, left hip, and the thorax joint. The z-axis is determined by the unit vector perpendicular to x-y plane and y-axis is the cross-product of x-axis and z-axis.
In the absence of ground-truth 3D poses in the target domain, we adopt the pseudo-labels generated in Section III-C as surrogate ground-truth labels that will be used to guide the global transformation process. Given an augmented 3D pose $\bm{P}_{3D}^{\mathcal{A}}$ from the source dataset, we aim to transform it using rotation and translation matrices to align with a 3D pseudo-pose $\tilde{\bm{P}}_{3D}^{tgt}$ sampled from the target dataset. First, $\bm{P}_{3D}^{\mathcal{A}}$ is centered at $H$, subtracting the root location $\bm{P}_{r}^{\mathcal{A}}$. Second, a rotation is applied to align the joints’ relative coordinates. Using the human coordinate system $H$ as a bridge, the rotation $\bm{R}$ from the source pose to the target pose can be decomposed to be (i) from the source camera coordinate system $C_{s}$ to $H$ and then (ii) from $H$ to the target camera coordinate system $C_{t}$. Third, a translation is applied to align the root of target pose $\tilde{\bm{P}}_{r}^{tgt}$. The entire transformation is
$\bm{P}_{3D}^{{}^{\prime}}=\bm{R}_{2}^{-1}\cdot\bm{R}_{1}\cdot(\bm{P}_{3D}^{\mathcal{A}}-\bm{P}_{r}^{\mathcal{A}})+\tilde{\bm{P}}_{r}^{tgt},$ (3)
where $\bm{R}_{1}$ and $\bm{R}_{2}$ denote the rotation matrices from $C_{s}$ to $H$ and $C_{t}$ to $H$ respectively, and ${\bm{P}}_{3D}^{{}^{\prime}}=[\bm{X}^{{}^{\prime}},\bm{Y}^{{}^{\prime}},\bm{Z}^{{}^{\prime}}]^{T}\in R^{3\times J}$ is the transformed 3D pose. See Appendix B for the details of computing the rotation matrix from a camera coordinate system to $H$.
@@
Finally, we can project ${\bm{P}}_{3D}^{{}^{\prime}}$ into 2D pose ${\bm{P}}_{2D}^{{}^{\prime}}=[\bm{u},\bm{v}]^{T}\in R^{2\times J}$ where $(\bm{u},\bm{v})$ denote the 2D image coordinates, obtained as
$\begin{array}[]{l}u_{j}=f_{x}\cdot\frac{X_{j}^{{}^{\prime}}}{Z_{j}^{{}^{\prime}}}+c_{x},\\[10.0pt] v_{j}=f_{y}\cdot\frac{Y_{j}^{{}^{\prime}}}{Z_{j}^{{}^{\prime}}}+c_{y},~j=\left\{0,...,J-1\right\},\end{array}$ (4)
@@
where $(f_{x},f_{y},c_{x},c_{y})$ are the target dataset camera’s intrinsic parameters. As shown in Fig. 5, after global transformation, 3D poses from the source and target domains are aligned. For 2D poses, their root positions are also aligned and the scales are similar. This alignment process significantly reduces the domain discrepancies caused by differences in camera viewpoints and positions. We analyse the effect of the proposed human-centered global transformation module compared with the Kabsch algorithm [50, 51] in Sec. IV-E.
@@
Figure 5: Visualization of the global transformation. A pseudo-3D pose is inferred from the target 2D pose. It is used to guide the rotation and translation of the source 3D pose for alignment. Given the aligned 3D poses, corresponding 2D poses are obtained through projection using the intrinsic camera parameters of the target dataset so that the 2D root positions are also aligned.
@@ III-F Iterative training architecture
After obtaining the aligned 2D and 3D poses ${\bm{P}}_{2D}^{{}^{\prime}}$ and ${\bm{P}}_{3D}^{{}^{\prime}}$ respectively, we train an adversarial pose augmentation framework to ensure the fidelity of the generated human poses (following previous works [17, 21]). This augmentation framework involves our generator, i.e., the pose augmentor and the discriminators $M_{d}^{2D}$, $M_{d}^{3D}$. The generator tries to generate fake samples as close as possible to real samples, while the discriminator tries to distinguish between them. $M_{d}^{2D}$ uses the aligned 2D pose ${\bm{P}}_{2D}^{{}^{\prime}}$ and the target 2D pose ${\bm{P}}_{2D}^{tgt}$ at the middle frame as input, to adversarially guide ${\bm{P}}_{2D}^{{}^{\prime}}$ to align with the target sample. This is unlike PoseAug [17] which uses the 2D poses of source dataset as input. The discriminator $M_{d}^{3D}$ uses the middle-frame 3D augmented pose ${\bm{P}}_{3D}^{\mathcal{A}}$ and the source pose ${\bm{P}}_{3D}^{src}$ as input to ensure the fidelity of poses after bone angle and bone length augmentation. The middle frame is taken as the input since nearby frames are expected to show minor variations, making the discriminator effective across the temporal window.
Inspired by Kinematic Chain Space (KCS) [54, 55], we use the KCS representation of 3D pose as input rather than 3D joints. KCS is a representation of human pose that captures the structural relationships between bones by computing the inner products of bone vectors. Instead of considering the entire body pose as a whole, we use a part-aware KCS [17] which separates the whole body into 5 parts: torso and left/right arm/leg. By focusing on local joint angles, this enlarges the feasible region of the augmented pose while maintaining plausibility and promoting diversity. To compute the part-aware KCS of an input pose $X$, we convert the pose to its bone direction vector $b$ first. Then, $b$ is separated corresponding to the five parts, denoted as $b_{i},i=1,...,5$. The local joint angle matrix $\mathcal{K}_{i}$ is calculated for each part as
$\mathcal{K}_{i}=b_{i}^{T}\cdot b_{i},$ (5)
which captures the inter-joint angle relationships within the $i$-th part. $\mathcal{K}_{i}$ is taken as the input to the 3D pose discriminator $M_{d}^{3D}$.
In each epoch, the generator $M_{a}$, discriminators $M_{d}^{2D}$, $M_{d}^{3D}$ and the pose estimator $M_{e}$ are trained jointly. $M_{e}$ is the pose estimation model that is fine-tuned to adapt to the target dataset. Their loss functions are as follows.
@@
Generator loss. The adversarial loss of the generator $M_{a}$ is the feedback from the discriminators:
@@
As shown in Eqs. 6 and 7, the generator is trained to minimize the discrepancy between the aligned 2D pose $\bm{P}_{2D}^{{}^{\prime}}$ and the target 2D pose $\bm{P}_{2D}^{tgt}$, as well as between the augmented 3D pose $\bm{P}_{3D}^{\mathcal{A}}$ and the source 3D pose $\bm{P}_{3D}^{src}$. In contrast, the discriminator aims to distinguish between these respective pose pairs. In this adversarial approach, the 3D pose discriminator $M^{3D}_{d}$ is used for evaluating the joint angle plausibility since the KCS representation encapsulates the inter-joint angle information. As the 2D poses contain information such as view point (rotation), position (translation), and body size (bone length), the 2D pose discriminator $M^{2D}_{d}$ is used for evaluating the body size, camera viewpoint, and position plausibility [17].
@@
where $\bm{P}$ and $\tilde{\bm{P}}$ are the ground truth and predicted 3D poses respectively. Before training $M_{e}$, the generator and the discriminators need to be warmed up for $m$ epoch.
After the training process, we obtain a fine-tuned Pose Estimator $M_{e}$ which generates predicted 3D poses $\hat{\bm{P}}_{3D}^{tgt}$ for the target dataset. These poses could be used as input for the Pseudo-Label Generation module $M_{p}$ again, and thus, the training process can be iterative and the number of iterations is studied in the ablation study.
@@ IV-A Datasets and Metrics
Human3.6M (H3.6M) is an extensive dataset for 3D human pose estimation, captured in a controlled indoor environment. The dataset includes over 3.6 million frames, covering seven main subjects (S1, S5, S6, S7, S8, S9, and S11) engaged in 15 diverse activities. Following previous works [21, 20], H3.6M has two main settings: (i) for cross-dataset evaluation, we use subjects S1, S5, S6, S7, and S8 as the source domain, with the other dataset as the target domain, and (ii) for cross-scenario evaluation, we use only S1 as the source domain, with S5, S6, S7, and S8 as the target domain. Measures MPJPE and PA-MPJPE are standard metrics employed to evaluate performance across these tasks.
@@ IV-B Implementation Details
Pose Estimator. Following previous works [17, 18, 22], we use single-frame-based VideoPose3D [23] and video-based MixSTE [24] for our Pose Estimator $M_{e}$. The input sequence length of MixSTE was set to 27 for 3DHP and 3DPW, and 243 for Human3.6M following the settings in [22]. The pre-trained weights from the source dataset are used as initial weights. We adopt the 16-keypoint human model with the 0-th joint (hip center point) as the origin and we use ground truth 2D keypoints as input.
Generators and discriminators. The generator $M_{a}$ has three 3-layer residual MLPs with LeakyReLU to generate augmented bone angles and bone lengths. We set $\alpha=1$ and $\beta=6$ in Eq. 7. The discriminators $M_{d}^{2D}$ and $M_{d}^{3D}$ use similar architectures for processing 2D and 3D poses. $M_{d}^{2D}$ incorporates 4-layer residual MLPs with LeakyReLU, taking the 2D keypoints as input. $M_{d}^{3D}$ consists of five 4-layer residual MLPs with LeakyReLU with each corresponding to a part of KCS representation.
Training. The initial learning rates for the discriminators, generator, and pose estimator are all set to 0.0001. They all use Adam optimizer. The experiments are conducted with a batch size of 1024 for 50 epochs and there is a warm-up phase lasting 2 epochs to train the generator and discriminators.
@@ IV-C Results
[…]
@@
[…]
@@
[…]
@@
Table III presents comparative results against PoseAug[17] and AdaptPose[18] with H3.6M as the source and 3DHP-C as the target dataset. We do not compare with PoseDA [21] as their complete codes are not available, nor with PoSynDA [22] since we obtained excessively incorrect results using their provided code. Our method outperforms by a large margin. This improvement stems from our explicit handling of camera viewpoint discrepancies which impresses the potential effectiveness of the proposed method under challenging viewpoints.
@@
Data Visualizations. Fig. 6 (a)-(d) presents example qualitative results on the 3DHP dataset which show that compared to the pre-trained baseline model VideoPose3D[23], our method performs significantly better, especially for joints that are farther away from the root joint. Fig. 6(e) illustrates a failure case of our method where the estimation of joints on the right side of the body is not accurate enough due to extreme self-occlusion.
Figure 6: Example 3D pose estimations on 3DHP dataset. (a)-(d) Our method (blue) performs significantly better than the pre-trained model (green) with smaller errors compared with the ground truth (red). (e) In cases of severe self-occlusion, our method fails to obtain sufficient estimations.
Next, we use UMAP [62] to perform a non-linear embedding into 2D and visualize the data distribution in source H3.6M, target 3DHP, and the generated 3D poses. As shown in Fig. 7, the generated 3D poses shows higher overlap with 3DHP than with H3.6M, indicating significant effectiveness and consistency with the target domain.
Figure 7: Distribution of 3D poses from H3.6M, 3DHP, and generated data. The 2D embeddings are produced with UMAP [62]. Our generated data is more distributionally aligned with the target domain (3DHP), exhibiting greater overlap.
In addition, we provide a more intuitive visualization of the coordinate distributions of selected distal joints after root subtraction. As shown in Fig. 8, the joint location distribution of the generated 3D poses closely matches that of 3DHP, demonstrating its effectiveness in simulating the target domain.
Figure 8: Distribution of the locations of joints from H3.6M, 3DHP, and generated 3D poses.
Furthermore, as shown in Fig. 9, we visualize the discrepancy in camera viewpoints by computing the histogram of camera viewpoint azimuth and elevation relative to the body-centered coordinate frame. The generated 3D poses mimic the camera viewpoint distribution of the target domain, further supporting its effectiveness in bridging the domain gap.
Figure 9: Distribution of camera viewpoint azimuth ($180^{\circ},180^{\circ}$) and elevation ($90^{\circ},90^{\circ}$) from H3.6M, 3DHP and generated 3D poses.